OpenStudy (anonymous):

Divide 20x^8 - 20x^4 + 12x^2 divide by 4x^2 =

OpenStudy (gw2011):

Divide each term by 4x^2 and you get: 20x^8/4x^2=5x^6 -20x^4/4x^2=-5x^2 12x^2/4x^2=3 Therefore, the answer is: 5x^6-5x^2+3